OpenStudy (anonymous):

Solve for (dy/dx) at the given points x^2-y^4=x^3+y^5 at (x,y) = (-1,1)

OpenStudy (anonymous):

2x -4y^3 dy/dx = 3x^2 +5y^4 dy/dx

OpenStudy (anonymous):

\[2x - 4y^3\frac{dy}{dx} = 3x^2 +5y^4 \frac{dy}{dx}\]

OpenStudy (anonymous):

u do the rest

OpenStudy (anonymous):

Many thanks elecengineer! Thats exactly the type of help I was looking for. I just wasn't very sure on how to to begin the problem.